OpenStudy (anonymous):

how to show the sequence {yn}, with yn=1/((2^2)^n) is at least superlinear ?

OpenStudy (anonymous):

for n=1 1^2=1*(1+1)(2*1+1)/6 1=1 let n=k 1^2+2^2+...+k^2=k(k+1)(2k+1)/6 we want to prove for n=k+1 1^2+2^2+...+k^2+(k+1)^2=(k+1)(k+2)(2k+… but 1^2+2^2+...+k^2=k(k+1)(2k+1)/6 then we must prove k(k+1)(2k+1)/6+(k+1)^2=(k+1)(k+2)(2k+3… divide by k+1 and maltiply by 6 k(2k+1)+6(k+1)=(k+2)(2k+3) 2k^2+k+6k+6=2k^2+3k+4k+6 2k^2+7k+6=2k^2+7k+6 true for all k

OpenStudy (anonymous):

replace all k by Y